OpenStudy (anonymous):

the first erm of a geometric series is 3, and the sum of the first term and the second term is 15. what is the sum of the first six term? A.1023 B.3906 C.4096 D.11718

OpenStudy (anonymous):

i know that the second term is 12

hartnn (hartnn):

the common ratio r= 2nd term / 1st term = 12/3 = ... ? now just use the formula for sum, with a1 = 3, and n = 6 \(\large S_n=a_1\dfrac{r^n-1}{r-1}\)

OpenStudy (anonymous):

i dont know what to plug in to what

hartnn (hartnn):

you could get r= 12/3 = 4, right ? plug in r= 4,a1 = 3 and n=6

OpenStudy (anonymous):

\[3\frac{ 95 }{3 }\]